OpenStudy (anonymous):

what is the perimeter of a triangle that has the sides of 6 and x+8 and hypotenuse of 6x

OpenStudy (anonymous):

\(6+x+8+6x=7x+14 \)?

OpenStudy (across):



OpenStudy (across):

Using the Pythagorean theorem, \[6^2+(x+8)^2=(6x)^2,\]\[36+x^2+16x+64=36x^2,\]\[-35x^2+16x+100=0,\]\(x\approx1.93426\) or \(x\approx-1.47712\). Since the second value of \(x\) makes no physical sense, use the first one by plugging it into FoolForMath's equation to calculate a perimeter.